What volume will be occupied by a balloon containing 0.65 mol H at 25 C and 1.0atm pressure ?
1
Expert's answer
2012-12-24T10:49:27-0500
H is incorrect ( H2 is correct)
We can use the next equation for solving this question

PV = nRT
n = 0.65 mol
T = 273 + 25 = 298 K
P = 1 atm = 101.1 kPa

V = nRT/P = 0.65*8.31*298/101.1 = 15.92 L